OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle (x + 6)2 + (y + 4)2 = 25 at the point (-9, -8).

OpenStudy (amistre64):

define the slope of the line that is between the center and the stated point

OpenStudy (amistre64):

then perp it by flipping and negating it a/b perps to -b/a

OpenStudy (amistre64):

then attach that perped slope to the stated point in your favorite line equation format

OpenStudy (anonymous):

Could you explain that a little differently?

OpenStudy (amistre64):

do you know how to define the slope between 2 points?

OpenStudy (anonymous):

I'm trying to recall but it's been awhile. can you remind me?

OpenStudy (amistre64):

change in y over change in x\[\frac{\Delta y}{\Delta x}=\frac{y_o-y_1}{x_o-x_1}\]

OpenStudy (anonymous):

Okay...

OpenStudy (amistre64):

-4--8 = 4 -6--9 = 3

OpenStudy (amistre64):

the tangent to a circle is 90 degrees (perpendicular) to the slope of the radius ...

OpenStudy (anonymous):

Right

OpenStudy (amistre64):

perp slopes have the property that their product is -1: \[\frac{4}{3}*\frac{y'}{x'}=-1\] \[\frac 34\frac{4}{3}*\frac{y'}{x'}=-1*\frac 34\] \[\frac{y'}{x'}=-\frac 34\]

OpenStudy (amistre64):

now, given a point (a,b) and a slope (m), we can define the equation of a line as:\[y=m(x-a)+b\]

OpenStudy (anonymous):

W

OpenStudy (anonymous):

Where did the 'a' come from?

OpenStudy (amistre64):

in order to make a setup as generic as possible; you use placeholder (letters, variables, etc) to establish a format with

OpenStudy (amistre64):

the generic point (a,b) is made specific by the specific point (-9,-8)

OpenStudy (anonymous):

Okay okay sorry I knew that

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

Thank you yet again